# DRAN: A Distribution and Relation Adaptive Network for Spatio-temporal Forecasting

### v1

@@ VII Conclusion
To conclude, Spatio-temporal forecasting is essential for understanding the states of complex systems, yet accurate predictions are often hindered by the dynamic and intricate nature of these systems. This study addresses the challenge of adapting to dynamic changes in spatio-temporal systems using neural networks. We propose a DRAN to accommodate changes in distribution shifts, relations, and stochastic variations. Our approach includes a SFL to enable effective temporal normalization for spatio-temporal contexts. Additionally, we develop a DSFL to capture features from both dynamic and static relations. Furthermore, our framework enables to learn the deterministic and stochastic representations of features. Experimental results demonstrate the superiority of our method and the effectiveness of its components.

### v2

@@ VII Conclusion
To conclude, spatio-temporal forecasting is essential for understanding the states of complex systems, yet accurate predictions are often hindered by the dynamic and intricate nature of these systems. This study addresses the challenge of adapting to dynamic changes in spatio-temporal systems using neural networks. A DRAN framework is proposed to accommodate changes in distribution shifts, relations, and stochastic variations. Our approach includes a SFL to enable effective temporal normalization for spatio-temporal contexts. Additionally, we develop a DSFL to capture features from both dynamic and static relations. Furthermore, our framework enables to learn the deterministic and stochastic representations of features. Experimental results demonstrate the superiority of our method and the effectiveness of its components.